A club consists of 25 married couples. Determine how many ways a committee of 7 people can be formed from these couples if
a) there are no restrictions?
b) there are exactly 5 men on the committe?
c) at least one person must be a man?
d)there must be at least 3 women and at least 2 men?
: A club consists of 25 married couples. Determine how many ways a committee of 7 people can be formed from these couples if
a) there are no restrictions?
b) there are exactly 5 men on the committe?
c) at least one person must be a man?
d)there must be at least 3 women and at least 2 men?

Answer by stanbon(19743)

A club consists of 25 married couples. Determine how many ways a committee of 7 people can be formed from these couples if
a) there are no restrictions?
Ans:
50C7 = 99,884,400
--------------------------------------------
b) there are exactly 5 men on the committee?
Ans:
25C5 * 25C2 = 53130*300 = 15,939,000
--------------------------------------------

c) at least one person must be a man?
# with NO men is 25C7 = 480,700
So, # with at least one man is 50C7-480,700 = 99,403,700
-----------------------------------------------------
d)there must be at least 3 women and at least 2 men?
That's pretty messy so I'll leave that to you.
=============================
Cheers,
Stan H.